OpenStudy (babyslapmafro):

When finding the derivative of ln|x| Why is the derivative 1/x and not 1 divided by the absolute value of x?

OpenStudy (amistre64):

\[D[ln(abs(x))]\] \[D[ln(sqrt(x^2))]*D[sqrt(...)]*D[x^2]\] \[\frac{1}{sqrt(x^2)}*\frac{1}{2sqrt(x^2)}*2x\] \[\frac{1}{|x|}*\frac{1}{2|x|}*2x\] \[\frac{x}{|x||x|}\] when x is a negative value, we still have a negative slope \[\frac{-3}{|-3||-3|}=-\frac{1}{3}\]